# Data Augmentation using LLMs: Data Perspectives, Learning Paradigms and Challenges

### v1

@@
Abstract
@@ 1 Introduction
Data-centric approaches to Artificial Intelligence (AI) constitute a pivotal element in the advancement towards Artificial General Intelligence (AGI), centering on the construction of AI systems underpinned by high-quality data Zha et al. (2023). This emphasis on data quality is vital, as it ensures the clarity of the information from which AI systems are to learn. Nevertheless, the acquisition of high-quality data presents significant challenges, being both costly and time-intensive, while the data annotation phase is often laborious and prone to inaccuracies resulting from human involvement Ding et al. (2022b). In response to these challenges, researchers have dedicated efforts towards data augmentation (DA) techniques as a means to mitigate such issues Chen et al. (2023b). Data augmentation fundamentally involves the adoption of innovative methods aimed at bolstering model efficacy through the broadening of training data diversity, all without necessitating further data collection efforts. This strategy effectively tackles a major hindrance in machine learning research, namely the dearth of readily accessible training data, by applying slight modifications to existing datasets or creating synthetic data instances.
@@
forked edges, for tree = grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em, font=, where level=2text width=7.0em, font=, where level=3text width=11.0em, font=, where level=4text width=8.0em, font=, where level=5text width=8.0em, font=, [ Data Perspectives, fill=lavender [ Data Perspectives, fill=lavender [ Data Creation [ e.g. (Chintagunta et al., 2021; Ho et al., 2022; Zheng et al., 2023a; Whitehouse et al., 2023; Shum et al., 2023), text width=33em, ] ] [ Data Labeling [ e.g. (Törnberg, 2023; Zhu et al., 2023b; Labruna et al., 2023; Zhang et al., 2023a; Meoni et al., 2023; Wang et al., 2023a), text width=33em, ] ] [ Data Reformation [ e.g. (Dixit et al., 2022; Korenčić et al., 2022; Chen et al., 2023c; Dunlap et al., 2023; Sharma et al., 2023a; Dai et al., 2023), text width=33em, ] ] [ Co-annotation [ e.g. (Li et al., 2023a; Bertaglia et al., 2023; Lu et al., 2023; Zhang et al., 2023d), text width=33em, ] ] ] [ Learning Paradigm, fill=lavender [Generative Learning [ Supervised Instruction Learning [ e.g. (Wang et al., 2023d; Taori et al., 2023; Peng et al., 2023), text width=33em, ] ] [ In-context Learning [ e.g. (Jiao et al., 2023; Wang et al., 2023c; Kong et al., 2024; Banerjee et al., 2024; Xiong et al., 2024; Luo et al., 2024a), text width=33em, ] ] [ Alignment Learning [ e.g. (Rafailov et al., 2023; Tang et al., 2023; Zhao et al., 2023a; Dong et al., 2023a; Wang et al., 2024), text width=33em, ] ] ] [Discriminative Learning [ Generating Pseudo Data for Classification [ e.g. (Wang et al., 2021; Ding et al., 2022b; Dai et al., 2023; He et al., 2023a; Belal et al., 2023; Fang et al., 2023), text width=33em, ] ] [ Scoring Data for Regression [ e.g. (Kwon et al., 2023; Klissarov et al., 2023; Ma et al., 2023; Liu et al., 2023f; Hasanbeig et al., 2023; Cao et al., 2024), text width=33em, ] ] ]] [ Challenges, fill=lavender [ Data Contamination in DA [ e.g. (Dodge et al., 2021; Magar and Schwartz, 2022; Sainz et al., 2023; Li et al., 2023e), text width=33em, ] ] [ Controllable DA [ e.g. (Liu et al., 2020; Chen et al., 2023a; Liu et al., 2023d, c; Yuan et al., 2023b), text width=33em, ] ] [ Culture-Aware DA [ e.g. (Ding et al., 2022a; Yao et al., 2023; Huang and Yang, 2023; Liu et al., 2023a), text width=33em, ] ] [ Multimodal DA [ e.g. (He et al., 2023b; Karjus, 2023; Zhang et al., 2023b; Brooks et al., 2024; Pan et al., 2024), text width=33em, ] ] ] [ Tasks (in Appendix), fill=palepink ] [ Domains (in Appendix), fill=palepink ] ]
Figure 1: Taxonomy of Data Augmentation using LLMs. Tasks and Domains are included in the Appendix.
@@ 2 Related Surveys
Prior to the advent of Large Language Models (LLMs), surveys like those by Feng et al. (2021) and Hedderich et al. (2021) explore data augmentation in NLP, highlighting methodologies, applications, and the challenges of low-resource NLP. However, the surveyed approaches are constrained by the need for a deep linguistic understanding of traditional NLP tasks and do not fully utilize LLMs. The emergence of LLMs has revolutionized this landscape by significantly enhancing the quality of synthetic data, leading to innovative methodologies and broader applications. LLMs’ nuanced understanding of language has mitigated previous limitations such as the issue of synthetic data being of poor quality, containing significant noise, and detrimentally affecting model performance. This advancement has made data augmentation more effective and accessible. Chen et al. (2023b) provide an empirical analysis of various augmentation techniques in supervised and semi-supervised contexts, focusing on pre-LLM methods. However, unlike previous surveys, we focus on discussing data augmentation using LLMs from the data and learning perspectives.
@@ 3.3 Data Reformation
Such reformation techniques could naturally aid in counterfactual generation tasks, which reforms existing data to its counterfactual version. Disco (Chen et al., 2023c) uses LLMs to generate high-quality counterfactual data at scale. It first uses in-context learning with GPT-3 to generate phrasal perturbations, then uses a task-specific teacher model to filter and distill high-quality counterfactual data. Models trained using the generated counterfactuals display improved robustness and generalization across distributions. For retrieval-augmented generation, CORE (Dixit et al., 2022) uses GPT-3 to generate counterfactual edits to the input conditioned on the retrieved excerpts. The perturbations then help mitigate model bias and improve performance on out-of-distribution (OOD) data.
@@ 4 Learning Paradigms
Figure 2: Illustrations of Teacher-Student Learning.
@@ 4.1 Generative Learning
: Supervised Instruction Learning, leveraging LLMs for data augmentation, presents a novel approach to enhancing instructional data generation. The generation process entails creating instructional data in the form of (instruction, x, y) triplets, where ’x’ represents optional context or input, and ’y’ denotes the output generated by LLMs in response to the given instruction. Wang et al. (2023d) first introduced "Self-Instruct," a method to enhance the instruction-following ability of pre-existing language models by leveraging their own outputs. Taori et al. (2023) created the Alpaca dataset at low cost using Supervised Instruction Learning and Peng et al. (2023) further explored the potential of leveraging LLMs to create cross-lingual instruction following data, self-evaluation and cross-model assessment data and responses to unnatural instructions.
@@
: In-context Learning requires LLMs to continuously improve their performance in real-time interactions, tailoring their output to specific prompts or contexts without the need for explicit fine-tuning on labeled data. To enable LLMs with such a capability, data with rich context should be provided to LLMs during the training process. Kong et al. (2024) use existing LLMs to generate two multiturn dialogue datasets based on detailed annotations, enabling model’s in-context dialogue capabilities. Banerjee et al. (2024) utilize KGs to build up in-context training samples for LLM instruction tuning. Xiong et al. (2024) construct a synthetic temporal QA dataset to enhance the LLM’s capability in temporal QA tasks. Luo et al. (2024a) design structure-based augmentation methods with LLMs for Temporal Knowledge Graph Completion. It is a common trend that various models adopt LLMs to generate synthetic data with long context to enhance the model’s in-context dialogue, inference, QA, and Natural Language Generation (NLG) capabilities Touvron et al. (2023); Jiao et al. (2023); Wang et al. (2023c).
@@ 5.2 Controllable Data Augmentation
Controllable data augmentation using LLMs faces significant challenges, primarily due to difficulties in ensuring the quality of generated synthetic data Liu et al. (2020); Chen et al. (2023a). These techniques, aimed at enriching few-shot or low-resource datasets by targeting specific attributes, rely on models understanding and applying attribute-value mappings. Current methods, including In-context Learning and fine-tuning, often struggle with maintaining quality across non-target dimensions during augmentation Liu et al. (2023d).
@@
To address these challenges, we need to explore new methods that can decompose and reconstruct texts with LLMs to precisely control attribute changes. Moreover, the generated synthetic data must be approached with a keen awareness of the potential for bias and the ethical implications associated with LLM outputs. By adopting responsible augmentation strategies that promote diversity and implement ethical guidelines, such as integrating the in-context vector into data augmentation framework Liu et al. (2023c), we can mitigate biases and advance NLP in a manner that is technologically effective and ethically responsible Liu et al. (2023c); Yuan et al. (2023b).

### v2

@@
Data Augmentation using LLMs:
Data Perspectives, Learning Paradigms and Challenges
Bosheng Ding  1  Chengwei Qin11footnotemark: 1  1  Ruochen Zhao11footnotemark: 1  1   Tianze Luo1    Xinze Li1
Guizhen Chen1   Wenhan Xia2
Junjie Hu3   Anh Tuan Luu1   Shafiq Joty1,4
1
Nanyang Technological University, Singapore 2Princeton University
3University of Wisconsin–Madison 4Salesforce AI
{bosheng001, chengwei003, ruochen002, tianze001, xinze002, guizhen001, anhtuan.luu}@ntu.edu.sg;
wxia@princeton.edu; junjie.hu@wisc.edu; sjoty@salesforce.com
  Equal contribution, order decided by coin flip.
Tóm tắt nội dung
@@ 1 Introduction
Data-centric approaches to Artificial Intelligence (AI) constitute a pivotal element in the advancement towards Artificial General Intelligence (AGI), centering on the construction of AI systems underpinned by high-quality data Zha et al. (2023). This emphasis on data quality is vital, as it ensures the clarity of the information from which AI systems are to learn. Nevertheless, the acquisition of high-quality data presents significant challenges, being both costly and time-intensive, while the data annotation phase is often laborious and prone to inaccuracies resulting from human involvement Ding et al. (2022b). In response to these challenges, researchers have dedicated efforts towards data augmentation (DA) techniques as a means to mitigate such issues Chen et al. (2023c). Data augmentation fundamentally involves the adoption of innovative methods aimed at bolstering model efficacy through the broadening of training data diversity, all without necessitating further data collection efforts. This strategy effectively tackles a major hindrance in machine learning research, namely the dearth of readily accessible training data, by applying slight modifications to existing datasets or creating synthetic data instances.
@@
forked edges, for tree = grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em, font=, where level=2text width=7.0em, font=, where level=3text width=11.0em, font=, where level=4text width=8.0em, font=, where level=5text width=8.0em, font=, [ Data Perspectives, fill=lavender [ Data Perspectives, fill=lavender [ Data Creation [ e.g. (Chintagunta et al., 2021; Ho et al., 2022; Zheng et al., 2023a; Whitehouse et al., 2023; Shum et al., 2023), text width=33em, ] ] [ Data Labeling [ e.g. (Törnberg, 2023; Zhu et al., 2023b; Labruna et al., 2023; Zhang et al., 2023a; Meoni et al., 2023; Wang et al., 2023a), text width=33em, ] ] [ Data Reformation [ e.g. (Dixit et al., 2022; Korenčić et al., 2022; Chen et al., 2023d; Dunlap et al., 2023; Sharma et al., 2023a; Dai et al., 2023), text width=33em, ] ] [ Co-annotation [ e.g. (Li et al., 2023a; Bertaglia et al., 2023; Lu et al., 2023; Zhang et al., 2023d), text width=33em, ] ] ] [ Learning Paradigm, fill=lavender [Generative Learning [ Supervised Instruction Learning [ e.g. (Wang et al., 2023d; Taori et al., 2023; Peng et al., 2023), text width=33em, ] ] [ In-context Learning [ e.g. (Jiao et al., 2023; Wang et al., 2023c; Kong et al., 2024; Banerjee et al., 2024; Xiong et al., 2024; Luo et al., 2024a), text width=33em, ] ] [ Alignment Learning [ e.g. (Rafailov et al., 2023; Tang et al., 2023; Zhao et al., 2023a; Dong et al., 2023a; Wang et al., 2024), text width=33em, ] ] ] [Discriminative Learning [ Generating Pseudo Data for Classification [ e.g. (Wang et al., 2021; Ding et al., 2022b; Dai et al., 2023; He et al., 2023a; Belal et al., 2023; Fang et al., 2023), text width=33em, ] ] [ Scoring Data for Regression [ e.g. (Kwon et al., 2023; Klissarov et al., 2023; Ma et al., 2023; Liu et al., 2023f; Hasanbeig et al., 2023; Cao et al., 2024), text width=33em, ] ] ]] [ Challenges, fill=lavender [ Data Contamination in DA [ e.g. (Dodge et al., 2021; Magar and Schwartz, 2022; Sainz et al., 2023; Li et al., 2023e), text width=33em, ] ] [ Controllable DA [ e.g. (Liu et al., 2020; Chen et al., 2023b; Liu et al., 2023d, c; Yuan et al., 2023b), text width=33em, ] ] [ Culture-Aware DA [ e.g. (Ding et al., 2022a; Yao et al., 2023; Huang and Yang, 2023; Liu et al., 2023a), text width=33em, ] ] [ Multimodal DA [ e.g. (He et al., 2023b; Karjus, 2023; Zhang et al., 2023b; Brooks et al., 2024; Pan et al., 2024), text width=33em, ] ] [ Privacy-Preserving DA [ e.g. (Cilliers, 2019; Song et al., 2024; Yuan et al., 2023a; Behnia et al., 2022; Chen et al., 2023a), text width=33em, ] ] ] [ Tasks (in Appendix), fill=palepink ] [ Domains (in Appendix), fill=palepink ] ]
Hình 1: Taxonomy of Data Augmentation using LLMs. Tasks and Domains are included in the Appendix.
@@ 2 Related Surveys
Prior to the advent of Large Language Models (LLMs), surveys like those by Feng et al. (2021) and Hedderich et al. (2021) explore data augmentation in NLP, highlighting methodologies, applications, and the challenges of low-resource NLP. However, the surveyed approaches are constrained by the need for a deep linguistic understanding of traditional NLP tasks and do not fully utilize LLMs. The emergence of LLMs has revolutionized this landscape by significantly enhancing the quality of synthetic data, leading to innovative methodologies and broader applications. LLMs’ nuanced understanding of language has mitigated previous limitations such as the issue of synthetic data being of poor quality, containing significant noise, and detrimentally affecting model performance. This advancement has made data augmentation more effective and accessible. Chen et al. (2023c) provide an empirical analysis of various augmentation techniques in supervised and semi-supervised contexts, focusing on pre-LLM methods. However, unlike previous surveys, we focus on discussing data augmentation using LLMs from the data and learning perspectives.
@@ 3.3 Data Reformation
Such reformation techniques could naturally aid in counterfactual generation tasks, which reforms existing data to its counterfactual version. Disco (Chen et al., 2023d) uses LLMs to generate high-quality counterfactual data at scale. It first uses in-context learning with GPT-3 to generate phrasal perturbations, then uses a task-specific teacher model to filter and distill high-quality counterfactual data. Models trained using the generated counterfactuals display improved robustness and generalization across distributions. For retrieval-augmented generation, CORE (Dixit et al., 2022) uses GPT-3 to generate counterfactual edits to the input conditioned on the retrieved excerpts. The perturbations then help mitigate model bias and improve performance on out-of-distribution (OOD) data.
@@ 4 Learning Paradigms
Hình 2: Illustrations of Teacher-Student Learning.
@@ 4.1 Generative Learning
: Supervised Instruction Learning, leveraging LLMs for data augmentation, presents a novel approach to enhancing instructional data generation. The generation process entails creating instructional data in the form of (instruction, x, y) triplets, where ’x’ represents optional context or input, and ’y’ denotes the output generated by LLMs in response to the given instruction. Wang et al. (2023d) first introduced ”Self-Instruct,”a method to enhance the instruction-following ability of pre-existing language models by leveraging their own outputs. Taori et al. (2023) created the Alpaca dataset at low cost using Supervised Instruction Learning and Peng et al. (2023) further explored the potential of leveraging LLMs to create cross-lingual instruction following data, self-evaluation and cross-model assessment data and responses to unnatural instructions.
@@
: In-context learning represents a new paradigm where models, particularly LLMs like GPT-3, generate predictions based on examples provided within the context of the input, without requiring explicit retraining or parameter updates Liu et al. (2022); Dong et al. (2023b); Zhang et al. (2024). To enable LLMs with such a capability, data with rich context should be provided to LLMs during the training process. Kong et al. (2024) use existing LLMs to generate two multiturn dialogue datasets based on detailed annotations, enabling model’s in-context dialogue capabilities. Banerjee et al. (2024) utilize KGs to build up in-context training samples for LLM instruction tuning. Xiong et al. (2024) construct a synthetic temporal QA dataset to enhance the LLM’s capability in temporal QA tasks. Luo et al. (2024a) design structure-based augmentation methods with LLMs for Temporal Knowledge Graph Completion. It is a common trend that various models adopt LLMs to generate synthetic data with long context to enhance the model’s in-context dialogue, inference, QA, and Natural Language Generation (NLG) capabilities Touvron et al. (2023); Jiao et al. (2023); Wang et al. (2023c).
@@ 5.2 Controllable Data Augmentation
Controllable data augmentation using LLMs faces significant challenges, primarily due to difficulties in ensuring the quality of generated synthetic data Liu et al. (2020); Chen et al. (2023b). These techniques, aimed at enriching few-shot or low-resource datasets by targeting specific attributes, rely on models understanding and applying attribute-value mappings. Current methods, including in-context learning and fine-tuning, often struggle with maintaining quality across non-target dimensions during augmentation Liu et al. (2023d). Furthermore, model collapse is a critical challenge that arises when models learn from data generated by other models, leading to a degeneration where models progressively lose information about the true underlying data distribution Shumailov et al. (2023). Over time, this process results in models that forget the true underlying data distribution and converge to a narrow interpretation of the data with very little variance, potentially compromising the diversity and richness of the model’s output.
@@
To address these challenges, we need to explore new methods that can decompose and reconstruct texts with LLMs to precisely control attribute changes. To mitigate the issues of model collapse, the generated data should follow the data distribution of human-generated data to preserve integrity and reliability. Moreover, the generated synthetic data must be approached with a keen awareness of the potential for bias and the ethical implications associated with LLM outputs. By adopting responsible augmentation strategies that promote diversity and implement ethical guidelines, such as integrating the in-context vector Liu et al. (2023c) into the data augmentation framework, we can mitigate biases and advance NLP in a manner that is technologically effective and ethically responsible Liu et al. (2023c); Yuan et al. (2023b).
@@ 5.4 Multimodal Data Augmentation
5.5 Privacy issues of Data Augmentation
Challenges
Data augmentation using LLMs presents significant privacy concerns, especially in sensitive sectors like healthcare Cilliers (2019), where confidentiality is paramount. While LLMs offer the potential to enrich sparse datasets with synthetic data, safeguarding sensitive information within these datasets poses considerable challenges. Research by Song et al. (2024) underscores the limitations of Differential privacy (DP) in effectively generating synthetic data without compromising privacy or data quality. Concurrently, Yuan et al. (2023a) explore LLMs’ utility in healthcare for patient-trial matching, proposing a privacy-aware approach that mitigates privacy risks by using desensitized data. These studies highlight the ongoing struggle to leverage LLMs for data augmentation while ensuring robust privacy protection, indicating a crucial area for further research and development in privacy-preserving techniques.
Future Directions
Future directions to mitigate these privacy issues may involve developing more robust anonymization techniques to ensure that no traceable data is reproduced in the outputs. Additionally, implementing differential privacy Behnia et al. (2022) and federated learning Chen et al. (2023a) could help minimize privacy risks by allowing data to be processed in a way that doesn’t reveal identifiable information. Transparent data usage policies and regular audits could also play a crucial role in maintaining user trust and ensuring compliance with evolving data protection regulations. These measures would be essential in balancing the benefits of LLMs in data augmentation with the need to protect individual privacy.